The first ionisation energy of potassium is 100kcal mol-1.Calculate the lowest possible frequency of light that can ionise a potassium atom.

Grade:12

1 Answers

vikas askiitian expert
13 years ago

from bohrs theory , E = Eo(1-1/n2)

 

E1 (ionisation energy) = 100Kcalmol-1 = 4.2*105joulesmol-1 

 

this E1 is for 1 mol , for 1 atom Eo will be E1/Na                        (Na = avagadro number)

 

Eo = E1/Na = 7*10-18 joules

 

 so , E = Eo(1-1/n2)

 

 E = hv                          (v is frequency of light emmited)

hv = 7*10-18 (1-1/n2)

 

h = 6.6*10-34 , n should be 2   for minimum frequency ...

 

hv = 7*10-18(1-1/4)

 v = 7.95*1015 hertz